# Gate-tunable nonlocal Josephson effect through magnetic van der Waals bilayers

### v1

@@ III Penetration of nonlocal correlations into nonsuperconducting regions
where we introduce Pauli matrices in spin, particle-hole and layer spaces: $\sigma_{k}$, $\tau_{k}$ and $\rho_{k}$ ($k=0,x,y,z$). $\hat{\mu}=\left(\begin{matrix}\mu_{1}&0\\ 0&\mu_{2}\\ \end{matrix}\right)$ is a matrix in layer space composed of on-site energies of the first and the second layers, respectively. $\bm{\hat{h}}=\left(\begin{matrix}\bm{h}_{1}&0\\ 0&\bm{h}_{2}\\ \end{matrix}\right)$ is also a matrix in the layer space containing the exchange field of both layers. Eq. (2) should be supplemented by the normalization condition
$\check{g}^{2}(\bm{R},\bm{p}_{F})=1.$ (3)
@@
In this work we only consider collinear magnetic orientations of $\bm{h}_{1}$ and $\bm{h}_{2}$. Then $\bm{F}$ is diagonal in spin space and takes the form $\bm{F}=\bm{F}_{\uparrow}(1+\sigma_{z})/2+\bm{F}_{\downarrow}(1-\sigma_{z})/2$. The anomalous Green’s functions $\bm{F}_{\uparrow}$ can be expressed as sum of four solutions physically corresponding to different types of Cooper pairs, which can exist in the bilayer, with different character of the penetration into the nonsuperconducting region:
@@
where $\Delta h=h_{1}-h_{2}$. $\bm{F}_{\downarrow}$ is obtained from $\bm{F}_{\uparrow}$ with the substitution $h_{1,2}\to-h_{1,2}$. Eigenvalues $\lambda_{k}$, which determine the oscillation periods of all 4 solutions, are presented in Fig. 2 as functions of $\Delta\mu=\mu_{1}-\mu_{2}$ for different types of magnetic bilayers: ferromagnetic [Figs. 2(a)-(c)], antiferromagnetic [(d)-(f)], bilayer consisting of a ferromagnetic and a paramagnetic layer [(g)-(i)] and a ferrimagnetic bilayer [(j)-(l)]. The difference of on-site energies of the monolayers $\Delta\mu$ can be varied by gating potential applied to one of the monolayers. Different horizontal rows of Fig. 2 correspond to increase of the interlayer coupling $t_{z}^{F}$ from top to bottom.
@@
The upper row in Fig. 2 representing the case of the lowest interlayer coupling can be easily interpreted in terms of intralayer (local) and interlayer (non-local) Cooper pairs, where the paired electrons reside at the same monolayer or at different monolayers, respectively, see Fig. 3 for sketch of such pairs. From Fig. 2(a) it is seen that $\lambda_{1}=\lambda_{2}=-2h/v_{F}$ (where $h=h_{1}=h_{2}$). The corresponding eigenvectors are $\hat{w}_{1}=(1,0,0,0)^{T}$ and $\hat{w}_{2}=(0,0,0,1)^{T}$ up to the zero order with respect to small parameter $t_{z}^{F}$. That is, these correlations are local correlations residing at the first and second monolayers, respectively. Physically they represent the well-known mesoscopic Fulde-Ferrel-Larkin-Ovchinnikov (FFLO) state. Such oscillations of the Cooper pairs in the ferromagnet were predicted long ago [4, 5] and lead to $0-\pi$ transitions in S/F/S JJs. However, two other proximity-induced pairing states corresponding to $\lambda_{3,4}=(-2h\pm\Delta\mu)/v_{F}$ and eigenvectors $\hat{w}_{3}=(0,1,0,0)^{T}$ and $\hat{w}_{4}=(0,0,1,0)^{T}$ represent non-local pairs consisting of electrons residing at different monolayers. Local pairs are not sensitive to $\Delta\mu$, what is natural because the paired electrons only feel the chemical potential of its own monolayer. Their oscillation period can depend on $\mu$ only via the dependence $v_{F}(\mu)$, but it requires much larger variations of the chemical potential than we study here. At the same time the behavior of non-local pairs is completely different. Their inverse oscillation period depends on $\Delta\mu$ linearly. The possibility to sustain such pairs, which oscillation period can be adjusted by gating, is a specific and very promising property of layered materials with a small number of atomic layers. Below we demonstrate how the gate-tunability of the non-local pairs opens perspectives of implementing the gate-controllable $0-\pi$ transitions in the appropriate JJs.
@@
As the interlayer hopping increases, local and non-local pairs hybridize and the eigenstates of the proximity pair correlations become their mixture. This case is shown in the middle row of Fig. 2. All four states are sensitive to gating because the eigenvectors of all the states have non-local components. With further increase of interlayer hopping, spatial behavior of the proximity-induced Cooper pairs evolves towards the case of isotropic 3D metal, when it is impossible to introduce the concept of local pair, since the size of Cooper pair in the direction of the $z$-axis significantly exceeds the interlayer distance. The case of the largest interlayer hopping considered by us is presented in the bottom row of Fig. 2. Although the corresponding value of the interlayer hopping is still significantly less than the intralayer hopping, i.e. the system under consideration is far from the case of the isotropic material, the evolution of spatial behavior of Cooper pairs is already clearly traced. Indeed, we can see that the dependence of the oscillation period of all Cooper pairs on $\Delta\mu$ becomes weaker and gradually disappears. The reason is that in the limit of large $t_{z}^{F}\gg(h_{1,2},\Delta\mu)$ two eigenstates correspond to conventional Cooper pairs with $\lambda_{1,2}\approx-(h_{1}+h_{2})/v_{F}$. The eigenvector of such pairs is $\hat{w}_{1,2}\approx(1,\pm 1,\pm 1,1)^{T}$, that is the wave functions of electrons making the pair are homogeneously distributed over both monolayers, forming a “normal pair” [see Fig. 3]. Physically it is obvious that electrons of this pair feel the on-site energies of both monolayers equally and for this reason it is insensitive to $\Delta\mu$, as it is well-known for such pairs in isotropic systems. The other two pairs have very rapidly oscillating wave function with $\lambda_{3,4}\approx\pm 2t_{z}^{F}/v_{F}$ and $\hat{w}_{3,4}=(-1,\pm 1,\mp 1,1)^{T}$.
@@ IV Josephson current through magnetic vdW bilayers
The current between sites $\bm{\eta}$ and $\bm{\eta}^{\prime}$ per unit length along the $y$-axis can be calculated via the solutions of the Bogolubov-de Gennes equation as follows:
@@ IV.1 Nonlocal contribution to Josephson current through AF bilayers
At first we present the results for the Josephson current via the AF bilayer with $h_{1}=-h_{2}=h$ assuming equal on-site energies of both magnetic layers (no gating). The critical Josephson current $j_{c}$ as a function of the length $L$ of the AF interlayer is demonstrated in Fig. 4. $j_{c}$ is defined as the value of the Josephson current $j(\chi)$ at $\chi>0$ with maximal amplitude, which can be positive (00-JJ) or negative ($\pi$-JJ). Here $\chi$ is the superconducting phase difference between the leads. Different curves correspond to different values of the interlayer hopping $t_{z}^{F}$. It is seen that at low $t_{z}^{F}$ the critical current manifests oscillations. At such low value of the interlayer hopping the Josephson current is mainly transferred by local Cooper pairs. The oscillations of the critical current originate from the oscillations of the local Cooper wave function in the ferromagnetic layers. If $\Delta\mu=0$ then the oscillation period of the local pairs is determined as $L_{l}=2\pi\xi_{F,l}$ with $\xi_{F,l}=v_{F}/2\sqrt{(h^{2}+t_{z}^{F2})}$. The Josephson current is carried by the pairs travelling along different trajectories. It is known that for ferromagnetic interlayers containing many atomic layers and at $\xi_{F}\ll L\ll\xi_{N}$ the oscillations of the critical current are superimposed by the power law decay $j_{c}\sim(L/\xi_{F})^{(1-n)/2}$, where $n$ is the dimensionality of the space [4, 51]. The decay is due to averaging of the contributions corresponding to different trajectories. The local pair contribution in our case behaves in the same way with the substitution $\xi_{F}\to\xi_{F,l}$. However, in the numerical calculations we do not restrict ourselves by the limit $L\ll\xi_{N}$ and for this reason this power law decay is further superimposed by the exponential decay $\sim e^{-L/\xi_{N}}$ [8].
@@
As the interlayer hopping increases, a non-oscillating contribution is added to the oscillating Josephson current, which decays at the normal coherence length and becomes dominant at large values of $t_{z}^{F}$. This contribution is due to the nonlocal pairs, which do not oscillate in the AF bilayer at $\Delta\mu=0$ according to the consideration presented in the previous section. The amplitude of such pairs contains an additional factor $t_{z}^{F}$ with respect to the amplitude of the local pairs and, therefore, their relative contribution to the Josephson current is $\sim t_{z}^{F2}$. Due to the growing contribution of the non-local pairs the $0-\pi$ transitions in the JJ disappear with increase of $t_{z}^{F}$. Although at first glance the considered system physically look similar (except for the thickness of the layers) to S/F/S JJs with a magnetic interlayer composed of two parallel thick ferromagnetic layers with opposite magnetizations [52], the predicted supercurrent behavior is fundamentally different. It was demonstrated that if the AF layer is composed of two parallel oppositely magnetized thick F layers, the interface between the layers (domain wall) provides an additional channel for the Josephson current with effectively reduced dimension [52]. In the ballistic case the additional contribution to the Josephson current was found to decay as $\delta j_{c}\sim(L/\xi_{F})^{(2-n)/2}$ for $n=2,3$. This reduced decay was superimposed by standard oscillation with the period $2\pi\xi_{F}$. In our case, the contribution of non-local pairs does not oscillate at all, since they represent a completely different physical phenomenon. Namely, they should be interpreted as pair correlations of electrons with opposite spins, belonging to different layers, and not as an additional channel for the Josephson current of reduced dimensionality, where, nevertheless, each of the paired electrons belongs to both layers simultaneously, which is the cause of oscillations.
@@
At the largest values of $t_{z}^{F}$ when the anisotropy between the intralayer and interlayer hoping already disappears (this case is represented by the curve corresponding to $t_{z}^{F}=1$ in Fig. 4), the difference between local and non-local pairs is erased, as discussed in the previous section. In this case, from the point of view of the proximity effect our layered antiferromagnetic metal is practically no different from a normal metal.
@@ IV.2 Gate-tunable $0-\pi$ transitions in JJs with AF interlayers
As it was demonstrated in Sec. III, oscillation period of the non-local Cooper pairs is sensitive to the difference $\Delta\mu$ between on-site energies of both layers, which the pair reside at. It is a specific property of the non-local pairs, which is not inherent to the local pairs and also disappears in the limit of isotropic metal, when wave functions of electrons are distributed over several layers. The pair oscillations manifest themselves as oscillations of the critical Josephson current. For this reason the sensitivity of the proximity pairs to $\Delta\mu$ results in the corresponding sensitivity of the critical current. Experimentally in vdW bilayers $\Delta\mu$ can be tuned by gating one of the layers. Possibility to use gating to tune on-site energy is an exciting property of vdW materials, which allow for high degree of controllability of superconductivity [53, 54, 55, 56, 57, 58], ferromagnetism [36, 59, 46] and also magnetic proximity effect in nonsuperconducting [60] and superconducting [61, 62, 63] vdW heterostructures. The dependence of the critical current on the gating potential can not only experimentally prove the presence of the non-local proximity-induced Cooper pairs, but is also of great interest for applications, as it results in the possibility of gate-tunable $0-\pi$ transitions, as demonstrated below.
@@
From Eq. (13) at $t_{z}^{F}\gg(\Delta\mu,h)$ and $h_{1}=-h_{2}=h$ one can obtain $\lambda_{1,2}\approx\mp h\Delta\mu/v_{F}t_{z}^{F}$ with $\hat{w}_{1,2}\approx(1,\pm 1,\pm 1,1)^{T}$ and $\lambda_{3,4}\approx\mp[2t_{z}^{F}+(h^{2}+\Delta\mu^{2}/4)/v_{F}t_{z}^{F}]$ with $\hat{w}_{3,4}\approx(-1,\mp 1,\pm 1,1)^{T}$. At $\Delta\mu=0$ the main contribution to the Josephson current is provided by the pairs corresponding to $\lambda_{1,2}$ with no spatial oscillations. The contribution of the pairs corresponding to $\lambda_{3,4}$ is also seen as small oscillations, accompanying the overall decay. The corresponding oscillation period, obtained from the results of the BdG calculations, is equal to $L_{BdG}=11a$, which is in a good agreement with the analytical formula $\frac{2\pi}{|\lambda_{3,4}|}=13a$. Yet the amplitudes of $\lambda_{3,4}$ pairs are small as compared to $\lambda_{1,2}$ pairs. The reason is that the eigenvectors $\hat{w}_{1,2}$ are even with respect to the interchange of both magnetic layers $1\leftrightarrow 2$ (even layer parity) and the eigenvectors $\hat{w}_{3,4}$ are odd with respect to such interchange (odd layer parity). The last type of pairs should not be generated at an interface with fully symmetric two-layer superconductor. We believe that in our case a small admixture of such pairs is generated at the interface with the superconductor due to the breaking of the layer symmetry because the superconducting leads are in direct contact only with the upper magnetic layer. At $\Delta\mu\neq 0$ the $\lambda_{1,2}$ pair become oscillating and due to this reason the corresponding current contribution manifests $0-\pi$ transition as a function of the interlayer length. The behavior of the current in Fig. 5 is in good agreement with the above approximate formulas. Namely, for $\Delta\mu=0.2,0.3,0.4t_{x}^{S}$ the periods $L_{BdG}\approx 172a,112a,85a$ are obtained from numerical results of BdG calculations, while the above formula gives $\frac{2\pi}{|\lambda_{1,2}|}=169a,112a,84a$. The small admixture of the $\lambda_{3,4}$ pairs is still seen in the curves corresponding to $\Delta\mu\neq 0$.
Figure 6: AF bilayer. Critical Josephson current as a function of the difference between on-site energies of magnetic layers $\Delta\mu$. $L=80a$, other parameters are the same as at Fig.5.
The strong sensitivity of oscillation period of the pairs to $\Delta\mu$ provides a possibility to implement $0-\pi$ transitions upon varying gating potential. It is demonstrated in Fig. 6, where the critical current is plotted as a function of $\Delta\mu$ for a given length of the AF interlayer.

### v2

@@ III.2 Proximity-induced superconducting correlations in a non-superconducting bilayer
In this work we only consider collinear magnetic orientations of $\bm{h}_{1}$ and $\bm{h}_{2}$. Then $\bm{F}$ is diagonal in spin space and takes the form $\bm{F}=\bm{F}_{\uparrow}(1+\sigma_{z})/2+\bm{F}_{\downarrow}(1-\sigma_{z})/2$. The anomalous Green’s functions $\bm{F}_{\uparrow}$ can be expressed as a sum of four solutions physically corresponding to different types of Cooper pairs, which can exist in the bilayer, with different character of the penetration into the nonsuperconducting region:
@@
where $\Delta h=h_{1}-h_{2}$. $\bm{F}_{\downarrow}$ is obtained from $\bm{F}_{\uparrow}$ with the substitution $h_{1,2}\to-h_{1,2}$. Eigenvalues $\lambda_{k}$, which determine the oscillation periods of all 4 solutions, are presented in Fig. 2 as functions of $\Delta\mu=\mu_{1}-\mu_{2}$ for different types of magnetic bilayers: ferromagnetic [Figs. 2(a)-(c)], antiferromagnetic [(d)-(f)], bilayer consisting of a ferromagnetic and a paramagnetic layer [(g)-(i)] and a ferrimagnetic bilayer [(j)-(l)]. The difference of on-site energies of the monolayers $\Delta\mu$ can be varied by the gating potential applied to one of the monolayers. The different horizontal rows of Fig. 2 correspond to an increase of the interlayer coupling $t_{z}^{F}$ from top to bottom.
@@
The upper row in Fig. 2, representing the case of the lowest interlayer coupling, can be easily interpreted in terms of intralayer (local) and interlayer (non-local) Cooper pairs, where the paired electrons reside at the same monolayer or at different monolayers, respectively, see Fig. 3 for sketch of such pairs. From Fig. 2(a) it is seen that $\lambda_{1}=\lambda_{2}=-2h/v_{F}$ (where $h=h_{1}=h_{2}$). The corresponding eigenvectors are $\hat{w}_{1}=(1,0,0,0)^{T}$ and $\hat{w}_{2}=(0,0,0,1)^{T}$ up to the zero order with respect to small parameter $t_{z}^{F}$. That is, these correlations are local correlations residing at the first and second monolayers, respectively. Physically they represent the well-known mesoscopic Fulde-Ferrel-Larkin-Ovchinnikov (FFLO) state. Such oscillations of the Cooper pairs in the ferromagnet were predicted long ago [6, 7] and lead to $0-\pi$ transitions in S/F/S JJs. However, two other proximity-induced pairing states corresponding to $\lambda_{3,4}=(-2h\pm\Delta\mu)/v_{F}$ and eigenvectors $\hat{w}_{3}=(0,1,0,0)^{T}$ and $\hat{w}_{4}=(0,0,1,0)^{T}$ represent non-local pairs consisting of electrons residing at different monolayers. Local pairs are not sensitive to $\Delta\mu$, which is natural because the paired electrons only feel the chemical potential of their own monolayer. Their oscillation period can depend on $\mu$ only via the dependence $v_{F}(\mu)$, but it requires much larger variations of the chemical potential to become pronounced here. At the same time the behavior of non-local pairs, which also represent finite-momentum FFLO-type pairing, is completely different. Their inverse oscillation period depends on $\Delta\mu$ linearly. The possibility to sustain such pairs, for which the oscillation period can be adjusted by gating, is a specific and very promising property of layered materials with a small number of atomic layers. Below we demonstrate how the gate-tunability of the non-local pairs opens perspectives of implementing the gate-controllable $0-\pi$ transitions in the appropriate JJs.
@@
As the interlayer hopping increases, local and non-local pairs hybridize and the eigenstates of the proximity pair correlations become their mixture. This case is shown in the middle row of Fig. 2. All four states are sensitive to gating because the eigenvectors of all the states have non-local components. With further increase of interlayer hopping, the spatial behavior of the proximity-induced Cooper pairs evolves towards the case of isotropic 3D metal, where it is impossible to introduce the concept of local pair, since the size of the Cooper pair in the direction of the $z$-axis significantly exceeds the interlayer distance. The case of the largest interlayer hopping considered is presented in the bottom row of Fig. 2. Although the corresponding value of the interlayer hopping is still significantly less than the intralayer hopping, i.e. the system under consideration is far from the case of the isotropic material, the evolution of spatial behavior of Cooper pairs is already clearly traced. Indeed, we can see that the dependence of the oscillation period of all Cooper pairs on $\Delta\mu$ becomes weaker and gradually disappears. The reason is that in the limit of large $t_{z}^{F}\gg(h_{1,2},\Delta\mu)$ two eigenstates correspond to conventional Cooper pairs with $\lambda_{1,2}\approx-(h_{1}+h_{2})/v_{F}$. The eigenvector of such pairs is $\hat{w}_{1,2}\approx(1,\pm 1,\pm 1,1)^{T}$, that is the wave functions of electrons making the pair are homogeneously distributed over both monolayers, forming a “normal pair” [see Fig. 3]. Physically it is obvious that electrons of this pair feel the on-site energies of both monolayers equally and for this reason it is insensitive to $\Delta\mu$, as it is well-known for such pairs in isotropic systems. The other two pairs have a very rapidly oscillating wave function with $\lambda_{3,4}\approx\pm 2t_{z}^{F}/v_{F}$ and $\hat{w}_{3,4}=(-1,\pm 1,\mp 1,1)^{T}$.
@@ IV Josephson current through magnetic vdW bilayers
The current between sites $\bm{\eta}$ and $\bm{\eta}^{\prime}$ per unit length along the $y$-axis can be calculated via the solutions of the Bogoliubov-de Gennes equation as follows:
@@ IV.1 Nonlocal contribution to Josephson current through AF bilayers
At first we present the results for the Josephson current via the AF bilayer with $h_{1}=-h_{2}=h$ assuming equal on-site energies of both magnetic layers (no gating). The critical Josephson current $j_{c}$ as a function of the length $L$ of the AF interlayer is demonstrated in Fig. 4. $j_{c}$ is defined as the value of the Josephson current $j(\chi)$ at $\chi>0$ with maximal amplitude, which can be positive (00-JJ) or negative ($\pi$-JJ). Here $\chi$ is the superconducting phase difference between the leads. Different curves correspond to different values of the interlayer hopping $t_{z}^{F}$. It is seen that at low $t_{z}^{F}$ the critical current manifests oscillations. At such low value of the interlayer hopping the Josephson current is mainly transferred by local Cooper pairs. The oscillations of the critical current originate from the oscillations of the local Cooper wave function in the ferromagnetic layers. If $\Delta\mu=0$ then the oscillation period of the local pairs is determined as $L_{l}=2\pi\xi_{F,l}$ with $\xi_{F,l}=v_{F}/2\sqrt{(h^{2}+t_{z}^{F2})}$. The Josephson current is carried by the pairs traveling along different trajectories. It is known that for ferromagnetic interlayers containing many atomic layers and at $\xi_{F}\ll L\ll\xi_{N}$ the oscillations of the critical current are superimposed by the power law decay $j_{c}\sim(L/\xi_{F})^{(1-n)/2}$, where $n$ is the dimensionality of the space [6, 67]. The decay is due to averaging of the contributions corresponding to different trajectories. The local pair contribution in our case behaves in the same way with the substitution $\xi_{F}\to\xi_{F,l}$. However, in the numerical calculations we do not restrict ourselves by the limit $L\ll\xi_{N}$ and for this reason this power law decay is further superimposed by the exponential decay $\sim e^{-L/\xi_{N}}$ [10].
@@
As the interlayer hopping increases, a non-oscillating contribution is added to the oscillating Josephson current, which decays at the normal coherence length and becomes dominant at large values of $t_{z}^{F}$. This contribution is due to the nonlocal pairs, which do not oscillate in the AF bilayer at $\Delta\mu=0$ according to the consideration presented in the previous section. The amplitude of such pairs contains an additional factor $t_{z}^{F}$ with respect to the amplitude of the local pairs and, therefore, their relative contribution to the Josephson current is $\sim t_{z}^{F2}$. Due to the growing contribution of the non-local pairs the $0-\pi$ transitions in the JJ disappear with increase of $t_{z}^{F}$. Although at first glance the considered system physically look similar (except for the thickness of the layers) to S/F/S JJs with a magnetic interlayer composed of two parallel thick ferromagnetic layers with opposite magnetizations [68], the predicted supercurrent behavior is fundamentally different. It was demonstrated that if the AF layer is composed of two parallel oppositely magnetized thick F layers, the interface between the layers (domain wall) provides an additional channel for the Josephson current with effectively reduced dimension [68]. In the ballistic case the additional contribution to the Josephson current was found to decay as $\delta j_{c}\sim(L/\xi_{F})^{(2-n)/2}$ for $n=2,3$. This reduced decay was superimposed by standard oscillation with the period $2\pi\xi_{F}$. In our case, the contribution of non-local pairs does not oscillate at all, since they represent a completely different physical phenomenon. Namely, they should be interpreted as pair correlations of electrons with opposite spins, belonging to different layers, and not as an additional channel for the Josephson current of reduced dimensionality, where, nevertheless, both paired electrons feel the same exchange field.
@@
At the largest values of $t_{z}^{F}$ when the anisotropy between the intralayer and interlayer hopping already disappears (this case is represented by the curve corresponding to $t_{z}^{F}=1$ in Fig. 4), the difference between local and non-local pairs is erased, as discussed in the previous section. In this case, from the point of view of the proximity effect our layered antiferromagnetic metal is practically no different from a normal metal.
@@ IV.2 Gate-tunable $0-\pi$ transitions in JJs with AF interlayers
As it was demonstrated in Sec. III, the oscillation period of the non-local Cooper pairs is sensitive to the difference $\Delta\mu$ between the on-site energies of both layers, in which the pair resides at. It is a specific property of the non-local pairs, which is not inherent to the local pairs and also disappears in the limit of an isotropic metal, when wave functions of the electrons are distributed over several layers. The pair oscillations manifest themselves as oscillations of the critical Josephson current. For this reason the sensitivity of the proximity pairs to $\Delta\mu$ results in the corresponding sensitivity of the critical current. Experimentally in vdW bilayers $\Delta\mu$ can be tuned by gating one of the layers. The possibility to use gating to tune on-site energy is an exciting property of vdW materials, which allows for a high degree of controllability of superconductivity [69, 70, 71, 72, 73, 74], ferromagnetism [38, 75, 62] and also magnetic proximity effect in nonsuperconducting [76] and superconducting [59, 60, 61] vdW heterostructures. The dependence of the critical current on the gating potential can not only experimentally prove the presence of the non-local proximity-induced Cooper pairs, but is also of great interest for applications, as it results in the possibility of gate-tunable $0-\pi$ transitions, as demonstrated below.
@@
From Eq. (29) at $t_{z}^{F}\gg(\Delta\mu,h)$ and $h_{1}=-h_{2}=h$ one can obtain $\lambda_{1,2}\approx\mp h\Delta\mu/v_{F}t_{z}^{F}$ with $\hat{w}_{1,2}\approx(1,\pm 1,\pm 1,1)^{T}$ and $\lambda_{3,4}\approx\mp[2t_{z}^{F}+(h^{2}+\Delta\mu^{2}/4)/v_{F}t_{z}^{F}]$ with $\hat{w}_{3,4}\approx(-1,\mp 1,\pm 1,1)^{T}$. At $\Delta\mu=0$ the main contribution to the Josephson current is provided by the pairs corresponding to $\lambda_{1,2}$ with no spatial oscillations. The contribution of the pairs corresponding to $\lambda_{3,4}$ is also seen as small oscillations, accompanying the overall decay. The corresponding oscillation period, obtained from the results of the BdG calculations, is equal to $L_{BdG}=11a$, which is in a good agreement with the analytical formula $\frac{2\pi}{|\lambda_{3,4}|}=13a$. Yet the amplitudes of $\lambda_{3,4}$ pairs are small as compared to $\lambda_{1,2}$ pairs. The reason is that the eigenvectors $\hat{w}_{1,2}$ are even with respect to the interchange of both magnetic layers $1\leftrightarrow 2$ (even layer parity) and the eigenvectors $\hat{w}_{3,4}$ are odd with respect to such interchange (odd layer parity). The last type of pairs should not be generated at an interface with a fully symmetric two-layer superconductor. We believe that in our case a small admixture of such pairs is generated at the interface with the superconductor due to the breaking of the layer symmetry because the superconducting leads are in direct contact only with the upper magnetic layer. At $\Delta\mu\neq 0$ the $\lambda_{1,2}$ pair become oscillating and for this reason the corresponding current contribution manifests a $0-\pi$ transition as a function of the interlayer length. The behavior of the current in Fig. 5 is in good agreement with the above approximate formulas. Namely, for $\Delta\mu=0.2,0.3,0.4t_{x}^{S}$ the periods $L_{BdG}\approx 172a,112a,85a$ are obtained from the numerical results of the BdG calculations, while the above formula gives $\frac{2\pi}{|\lambda_{1,2}|}=169a,112a,84a$. The small admixture of the $\lambda_{3,4}$ pairs is still seen in the curves corresponding to $\Delta\mu\neq 0$.
Figure 6: AF bilayer. Critical Josephson current as a function of the difference between on-site energies of magnetic layers $\Delta\mu$. $L=80a$, other parameters are the same as in Fig.5.
The strong sensitivity of the oscillation period of the pairs to $\Delta\mu$ provides a possibility to implement $0-\pi$ transitions upon varying the gating potential. It is demonstrated in Fig. 6, where the critical current is plotted as a function of $\Delta\mu$ for a given length of the AF interlayer.